OpenStudy (anonymous):

Twice a number increased by 12 is equal to 31 less than three times the number. Find the number

OpenStudy (dangerousjesse):

\(\normalsize\color{black}{12x+12=3x-31}\)

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

i thought it was 2x+12=31-3x

OpenStudy (dangerousjesse):

\(\normalsize\color{black}{2x+12=3x-31}\) \(\LARGE\color{white}{ \rm │ }\) I mean

OpenStudy (anonymous):

its 3x-31 or 31-3x?

OpenStudy (dangerousjesse):

\(\normalsize\color{black}{31}\) less than \(\normalsize\color{black}{3x}\) is \(\normalsize\color{black}{3x-31}\)

OpenStudy (anonymous):

oh ok can you help with another question?

OpenStudy (dangerousjesse):

Sure.

OpenStudy (anonymous):

im going to open another post so you could get a medal :)